# StoneDetector: Conventional and versatile code clone detection for Java

### v1

@@ 5 Related Work
Techniques and tools for clone detection address the problem of finding code fragments that share a certain degree of syntactic and/or semantic similarity. Approaches for the detection of code clones can be classified according to the used detection methodology [6]. Following a recent survey on this research topic by Nasrabadi et al. [6], we differentiate text-, token-, tree-, graph-, and metric-based methods as well as hybrid, Deep Learning, and dynamic approaches. For an extensive discussion of code clone techniques and tools, we refer the interested reader to one of the various surveys available on the topic [6, 7, 1, 8] and here provide a condensed overview of the most formative and relevant work.
@@ 5.1 Text-based Techniques
The raw source code is used as input for the first group of methods. The source code is either used directly as a sequence of strings for finding matching parts of two code fragments, or only very limited pre-processing is applied beforehand, e.g., removal of whitespace and/or code normalization. Despite their simplicity and limited effectiveness on Type 3/4 code clones, text-based methods are still used today due to their efficiency and easy use for different programming languages. Examples include the general-purpose clone detectors like Simian [^17] https://simian.quandarypeak.com/ or plagiarism checkers like MOSS [48] [^18] https://theory.stanford.edu/~aiken/moss/.
As one of the first approaches, Baker proposed the tool Dup [49]. The tool combines a consistent identifier/literal renaming scheme with suffix trees on hashed code lines for finding parametrized code matches, i.e., near-miss code clones. Similarly, Johnson proposed to use the Karp-Rabin algorithm and sliding windows for hashed code lines, called fingerprints, to find duplicated code fragments [50]. In order to also detect Type 2 clones, he removes whitespace and comments and maps all identifiers, keywords, and literals to a common placeholder. The text-based clone detector Vuddy uses whole-function hashes on pre-processed source code for generating function fingerprints. These fingerprints are then used as indexes for the implementation of efficient lookups for vulnerable code in large codebases [51]. A well-known and up-to-date clone detector is NiCad [9, 52], which is mostly text-based but also applies lightweight parsing for normalizing the raw source code so that noise is removed and code formats are standardized. The code pre-processing can be adjusted to different programming languages using a customizable grammar definition. NiCad is using the Longest Common Subsequence algorithm for finding matching code and is one of the clone detectors used in our comparative evaluation in Sect. 4. The tool identifies almost all Type 1/2 clones and a substantial part of Type 3 code clones [12]. Various other authors have used NiCad as a starting point for their own research, e.g., Ragkhitwetsagul and Krinke [53] applied NiCad to compiled and subsequently decompiled code for further code normalization, while Chen et al. [54] used NiCad on decompiled Bytecode to identify Android malware.
@@ 5.2 Token-based Techniques
CCFinder [57] lexes two source code fragments into token sequences, thereby transforming some tokens according to certain rules, e.g., replacing mapping identifiers to a common token. The thus normalized token sequences are then handed over to the suffix tree algorithm for finding matching code fragments. Note that, as some of the above-mentioned text-based tools work quite similarly, the differentiation between text- and token-based approaches to clone detection is rather not strict. In subsequent work, the authors provided a flexible tokenization mechanism for easily extending the tool towards additional programming languages [58]. The suffix tree in its generalized form is also used by Göde and Koschke in their tool iClones for incremental code clone detection, i.e., detection of code clones in evolving code bases [17]. CP-Miner uses frequent subsequence mining for finding interleaved token subsequences in case of reordered statements or inserted code within code lines [36]. The token-based clone detector SourcererCC uses optimized indexing and filtering techniques for reducing the number of candidate code fragments when searching for code clones. In this way, the tool supports scalable clone detection for large code repositories with hundreds of millions of lines of code [18]. Scaling clone detectors for large codebases has also been the focus of subsequent research [59]. Another token-based technique is described by Ragkhitwetsagul and Krinke [60], their tool Siamese uses token sequences with different levels of abstractions and methods from document retrieval on their N-gram vector representations. As a result, they are able to provide a scalable method for the detection of Type 1, 2, and 3 code clones.
While tools like NiCad and SourcererCC are quite good at finding exact and near-miss clones, Type 3 and 4 clones with lower syntactical similarity provide a challenge. More recent token-based clone detectors, like CCAligner [10], LVMapper [56], and NIL [20], therefore focus on code clones with many consecutive code edits or many code modifications scattered around the code. These so-called large-gap clone detectors often apply more sophisticated string metrics for aligning and matching token sequences, e.g., e-edit distance [10] or Smith-Waterman sequence alignment [55]. We include the tools SourcererCC, iClones, and CCAligner as representative token-based clone detectors in our evaluation in Sect. 4
@@ 5.4 Metric-based Techniques
5.5 Deep Learning Techniques
@@
In the last decade, the application of machine learning and Deep Learning has become more widespread in clone detection. Common to these approaches is the embedding of code fragments into a manually crafted or automatically learned and often high-dimensional vector space, wherein statistical relationships are used for predicting similar code. The ability to thus automatically learn and recognize the hidden patterns of code clones from, e.g., today’s large and omnipresent code repositories, should guarantee promising results. While in particular Deep Learning clone detectors have shown very good performance in the detection of Type 1 to Type 4 code clones lately, research has recently also highlighted common problems and pitfalls in the application of Deep Learning to clone detection, i.e., training/evaluation bias, overfitting, and missing ground truth [78, 79, 80].
@@
Prior approaches considered conventional machine learning methods. For instance, Cesare et al. use algorithms like naive Bayes and random forests for detecting package-level clones [81]. Clustering via k-Means can also be used for unsupervised learning of similar code and for finding code clones with respect to various similarity thresholds [82]. More recent work, however, has shifted towards the application of deep neural networks. In pioneering research, White et al. use an autoencoder architecture for unsupervised learning of abstract syntax trees and compute code similarities on the learned representation [83]. The Deep Learning clone detector CCLearner uses the frequencies of token categories in the abstract syntax trees of code fragments as features and learns to discriminate code clones from non-code clones using these features via supervised training of a deep neural network [84]. CDLH [85] is using the long short-term memory (LSTM) architecture for representation learning on abstract syntax trees and is trained to embed the resulting representation to hash values, such that code clones are close in terms of their Hamming distance. Similarly, Zhang et al. are using, in their architecture ASTNN, bidirectional recurrent units on embedded statement trees for encoding abstract syntax trees and identifying code clones [86]. Different settings for embeddings of abstract syntax trees are evaluated by Büch and Andrzejak [87], where they use a Siamese architecture that optimizes the embeddings of two abstract syntax trees such that the cosine similarity in the case of code clones is maximized. As an interesting result, pre-trained embeddings like Word2vec, Code2vec [24] have proven useful. Another lightweight approach for embedding abstract syntax trees is presented by Gao et al. [88], where the authors use random walks on the abstract syntax trees and embed the resulting token sequences similar to Word2vec. The resulting encodings of code fragments can then be compared via their Euclidean distances.
@@ 5.6 Dynamic Techniques
Dynamic methods focus on the computations of code instead of its syntax and usually target binary code. Therefore, dynamic analyses or testing is used to execute code fragments with sample inputs and to track their behavior. The resulting data can then be used to match semantically similar code and, in this way, in particular, to detect Type 4 code clones. As for all dynamic methods, scalability and execution time can be problematic, and accuracy is tightly coupled to the coverage of test inputs.
@@ 5.7 Hybrid Techniques
Token-based and tree- or graph-based approaches to clone detection are often used in such a way that a token-based technique is used for filtering code clone candidates which are afterward validated using the tree-/graph-based technique [107]. A combination of token- and tree-based clone detection is also used by Vislavski et al. [108] in their tool LICCA to detect cross-language code clones and by Wang et al. [109] in their tool CCStokener to speed up detection efficiency while retaining detection performance. Similar two-step pipelines have been presented for metric-based approaches, where the metric-based technique is used for prefiltering clone candidates and a graph- or text-based technique is used for validation [110, 111].
Further research focused on the role of different intermediate representations for code clone detection, e.g., considering Java source code and Bytecode. Selim et al. apply the clone detector CCFinder [57] on the Jimple representation of Java Bytecode and find source code clones that are not detected otherwise. They thus propose to use their approach to complement clone detectors working solely on the source code [112]. A similar method has been described for the usage of NiCad [9] on LLVM-based intermediate representations for C [113]. A more comprehensive approach, including a study of token-, tree-, and graph- based code representations and a combination thereof within the tool TACC is presented by Wang et al. [114], where they demonstrate an increased performance for the combined approach for all types of code clones. The usage of different code representations in combination with Deep Learning is also analyzed by Tufano et al. [94].
@@ 6 Conclusion
The method for clone detection has been implemented in the publicly available StoneDetector platform, which provides a versatile system for finding code clones in both Java source code and Bytecode, as well as supporting manifold experiments on clone detection’s parameters. Using the implementation in StoneDetector, we conducted a thorough evaluation of our approach with respect to the state-of-the-art benchmarks for Java code clone detection, i.e., BigCloneBench, Google Code Jam, GPTCloneBench, and Project CodeNet. Furthermore, we experimentally compared the performance of the code clone detection approach in StoneDetector to other state-of-the-art clone detectors, including NiCad, iClones, CCAligner, SourcererCC, Deckard, Oreo, and NIL. The experiments proved StoneDetector’s competitive capabilities in finding exact near-miss clones with respect to precision and recall. In addition, StoneDetector excels at identifying code clones with larger syntactical variance at high precision when compared to the other state-of-the-art tools. Further experiments demonstrated StoneDetector’s ability to analyze large code bases with up to multiple million lines of code. Besides its performance, we highlight the versatility of the platform for experimenting with different parameters for code clone detection, including the similarity threshold, minimal size of code clones, string metrics, e.g., Hamming distance, Longest Common Subsequence, and hashing methods, e.g., Locality-Sensitive Hashing. Eventually, StoneDetector’s ready-made extensibility to other programming languages and program representations is presented, as is shown for stack-based and register-based Bytecode formats.
While this paper focuses on conventional code clone detection and therefore omits an in-depth discussion and comparative evaluation with Deep Learning approaches to code clone detection, we want to analyze and incorporate such models into the StoneDetector platform in our future work. While we acknowledge the promising results of, in particular, large language models in recent work [97], as discussed in the related work section, we also assume that conventional methods of clone detection conjoined with their advantages, as presented in this paper, will play a role in the future. We though concede that the assessment of the advantages and disadvantages of conventional code clone detection in comparison with Deep Learning approaches, as well as their incorporation into a possible combined approach, requires further research and analysis. Besides, we are also interested in extending the StoneDetector platform with respect to other programming languages, which have not been in the research focus for code clone detection.

### v2

@@ 5 Related Work
Techniques and tools for clone detection address the problem of finding code fragments that share a certain degree of syntactic and/or semantic similarity. Approaches for the detection of code clones can be classified according to the used detection methodology [70]. Following a recent survey on this research topic by Nasrabadi et al. [70], we differentiate text-, token-, tree-, graph-, and metric-based methods as well as hybrid, Machine/Deep Learning, and dynamic approaches. For an extensive discussion of code clone techniques and tools, we refer the interested reader to one of the various surveys available on the topic [83, 70, 81, 2] and here provide a condensed overview of the most formative and relevant work.
@@ 5.1 Text-based Techniques
The raw source code is used as input for the first group of methods. The source code is either used directly as a sequence of strings for finding matching parts of two code fragments, or only very limited pre-processing is applied beforehand, e.g., removal of whitespace and/or code normalization. Despite their simplicity and limited effectiveness on Type 3/4 code clones, text-based methods are still used today due to their efficiency and easy use for different programming languages. Examples include the general-purpose clone detectors like Simian [^19] https://simian.quandarypeak.com/ or plagiarism checkers like MOSS [96] [^20] https://theory.stanford.edu/~aiken/moss/. As one of the first approaches, Baker proposed the tool Dup [10]. The tool combines a consistent identifier/literal renaming scheme with suffix trees on hashed code lines for finding parametrized code matches, i.e., near-miss code clones. Similarly, Johnson proposed to use the Karp-Rabin algorithm and sliding windows for hashed code lines, called fingerprints, to find duplicated code fragments [42]. In order to also detect Type 2 clones, he also removes whitespace and comments and maps all identifiers, keywords, and literals to a common placeholder. The text-based clone detector Vuddy uses whole-function hashes on pre-processed source code for generating function fingerprints. These fingerprints are then used as indexes for the implementation of efficient lookups for vulnerable code in large code bases [49]. A well-known and up-to-date clone detector is NiCad [85, 66], which is mostly text-based but also applies lightweight parsing for normalizing the raw source code so that noise is removed and code formats are standardized. The code pre-processing can be adjusted to different programming languages using a customizable grammar definition. NiCad is using the longest common subsequence algorithm for finding matching code and is also one of the clone detectors used in our comparative evaluation in Sect. 4. The tool identifies almost all Type 1/2 clones and a substantial part of Type 3 code clones [102]. Various other authors have used NiCad as a starting point for their own research, e.g., Ragkhitwetsagul and Krinke [79] applied NiCad to compiled and subsequently decompiled code for further code normalization, while Chen et al. [18] used NiCad on decompiled Bytecode to identify Android malware.
@@ 5.2 Token-based Techniques
CCFinder [44] lexes two source code fragments into token sequences, thereby transforming some tokens according to some rules, e.g., replacing mapping identifiers to a common token. The thus normalized token sequences are then handed over to the suffix tree algorithm for finding matching code fragments. Note that, as some of the above-mentioned text-based tools work quite similarly, the differentiation between text- and token-based approaches to clone detection is rather not strict. In subsequent work, the authors provided a flexible tokenization mechanism for easily extending the tool towards additional programming languages [98]. The suffix tree in its generalized form is also used by Göde and Koschke in their tool iClones for incremental code clone detection, i.e., detection of code clones in evolving code bases [31]. CP-Miner uses frequent subsequence mining for finding interleaved token subsequences in case of reordered statements or inserted code within code lines [61]. The token-based clone detector SourcererCC uses optimized indexing and filtering techniques for reducing the number of candidate code fragments when searching for code clones. In this way, the tool supports scalable clone detection for large code repositories with hundreds of millions of lines of code [88]. Scaling clone detectors for large codebases has also been the focus of subsequent research [57]. CloneWorks also uses a token-based approach combined with several techniques to scale and fasten clone detection, including Jaccard similarity metric, sub-block filtering heuristics, index creation and partitioning [104]. Another token-based technique is described by Ragkhitwetsagul and Krinke [80], their tool Siamese uses token sequences with different levels of abstractions and methods from document retrieval on their n-gram vector representations. As a result, they are able to provide a scalable method for the detection of Type 1, 2, and 3 code clones.
While tools like NiCad and SourcererCC are quite good at finding exact and near-miss clones, Type 3 and 4 clones with lower syntactical similarity provide a challenge. More recent token-based clone detectors, like CCAligner [110], LVMapper [117], and NIL [69], therefore focus on code clones with many consecutive code edits or many code modifications scattered around the code. These so-called large-gap clone detectors often apply more sophisticated string metrics for aligning and matching token sequences, e.g., e-edit distance [110] or Smith-Waterman sequence alignment [63]. We include the tools SourcererCC, iClones, CloneWorks, NIL, and CCAligner as representative token-based clone detectors in our evaluation in Sect. 4
@@ 5.4 Metric-based Techniques
5.5 Machine/Deep Learning Techniques
@@
In the last decade, the application of machine learning and Deep Learning has become more widespread in clone detection. Common to these approaches is the embedding of code fragments into a manually crafted or automatically learned and often high-dimensional vector space, wherein statistical relationships are used for predicting similar code. The ability to thus automatically learn and recognize the hidden patterns of code clones from, e.g., today’s large and omnipresent code repositories, should guarantee promising results. While in particular Deep Learning clone detectors have shown very good performance in the detection of Type 1 to Type 4 code clones lately, research has recently also highlighted common problems and pitfalls in the application of Deep Learning to clone detection, i.e., training/evaluation bias, overfitting, and missing ground truth [53, 92, 52, 21].
@@
Prior approaches considered conventional machine learning methods. For instance, Cesare et al. use algorithms like naive Bayes and random forests for detecting package-level clones [16]. Clustering via k-Means can also be used for unsupervised learning of similar code and for finding code clones with respect to various similarity thresholds [46]. More recent work, however, has shifted towards the application of deep neural networks. In pioneering research, White et al. use an autoencoder architecture for unsupervised learning of abstract syntax trees and compute code similarities on the learned representation [116]. The Deep Learning clone detector CCLearner uses the frequencies of token categories in the abstract syntax trees of code fragments as features and learns to discriminate code clones from non-code clones using these features via supervised training of a deep neural network [59]. CDLH [115] is using the long short-term memory (LSTM) architecture for representation learning on abstract syntax trees and is trained to embed the resulting representation to hash values, such that code clones are close in terms of their Hamming distance. Similarly, Zhang et al. are using in their architecture ASTNN bidirectional recurrent units on embedded statement trees for encoding abstract syntax trees and identifying code clones [121]. Different settings for embeddings of abstract syntax trees are evaluated by Büch and Andrzejak [13], where they use a siamese architecture that optimizes the embeddings of two abstract syntax trees such that the cosine similarity in the case of code clones is maximized. As an interesting result, pre-trained embeddings like Word2vec, Code2vec [6] have proven useful. Another lightweight approach for embedding abstract syntax trees is presented by Gao et al. [30], where the authors use random walks on the abstract syntax trees and embed the resulting token sequences similar to Word2vec. The resulting encodings of code fragments can then be compared via their Euclidean distances. Hu et al. [37] choose manual feature engineering to encode abstract syntax trees using different centrality values and then employ decision trees for finding matching code clones based on the resulting vector representations. Different similarity measures based upon code fragments’ token sequences are considered by Feng et al. [28], which feed the resulting feature vectors into several classical machine learning algorithms for detecting code clones.
@@ 5.6 Dynamic Techniques
Dynamic methods focus on the computations of code instead of its syntax and usually target binary code. Therefore, dynamic analyses or testing is used to execute code fragments with sample inputs and track their behavior. The resulting data can then be used to match semantically similar code and, in this way, in particular, to detect Type 4 code clones. As for all dynamic methods, scalability and execution time can be problematic, and accuracy is tightly coupled to the coverage of test inputs.
@@ 5.7 Hybrid Techniques
Token-based and tree- or graph-based approaches to clone detection are often used in such a way that a token-based technique is used for filtering code clone candidates which are afterward validated using the tree-/graph-based technique [19]. A combination of token- and tree-based clone detection is also used by Vislavski et al. [108] in their tool LICCA to detect cross-language code clones and by Wang et al. [113] in their tool CCStokener to speed up detection efficiency while retaining detection performance. Similar two-step pipelines have been presented for metric-based approaches, where the metric-based technique is used for prefiltering clone candidates and a graph- or text-based technique is used for validation [111, 50]. Further research focused on the role of different intermediate representations for code clone detection, e.g., considering Java source code and Bytecode. Selim et al. apply the clone detector CCFinder [44] on the Jimple representation of Java Bytecode and find source code clones that are not detected otherwise. They thus propose to use their approach to complement clone detectors working solely on the source code [97]. A similar method has been described for the usage of NiCad [85] on LLVM-based intermediate representations for C [15]. A more comprehensive approach, including a study of token-, tree-, and graph- based code representations and a combination thereof within the tool TACC is presented by Wang et al. [114], where they demonstrate an increased performance for the combined approach for all types of code clones. The usage of different code representations in combination with Deep Learning is also analyzed by Tufano et al. [106].
@@ 6 Conclusion
The method for clone detection has been implemented in the publicly available StoneDetector tool, which provides a versatile system for finding code clones in both Java source code and Bytecode, as well as supporting manifold experiments on clone detection’s parameters. Using the implementation in StoneDetector, we conducted a thorough evaluation of our approach with respect to the state-of-the-art benchmarks for Java code clone detection, i.e., BigCloneBench, Google Code Jam, GPTCloneBench, and Project CodeNet. Furthermore, we experimentally compared the performance of the code clone detection approach in StoneDetector to other state-of-the-art clone detectors, including NiCad, iClones, CCAligner, SourcererCC, Deckard, Oreo, CloneWorks and NIL. The experiments proved StoneDetector’s competitive capabilities in finding exact near-miss clones with respect to precision and recall. In addition, StoneDetector excels at identifying code clones with larger syntactical variance at high precision when compared to the other state-of-the-art tools. Further experiments demonstrated StoneDetector’s ability to analyze large code bases with up to multiple million lines of code. Besides its performance, we highlight the versatility of StoneDetector experimenting with different parameters for code clone detection, including the similarity threshold, minimal size of code clones, string metrics, e.g., Hamming distance, Longest Common Subsequence, and hashing methods, e.g., Locality-Sensitive Hashing. Eventually, StoneDetector’s ready-made extensibility to other programming languages and program representations is presented, as is shown for stack-based and register-based Bytecode formats.
While this paper focuses on conventional code clone detection and therefore omits an in-depth discussion and comparative evaluation with Deep Learning approaches to code clone detection, we want to analyze and incorporate such models into StoneDetector in our future work. While we acknowledge the promising results of, in particular, large language models in recent work [25], as discussed in the related work section, we also assume that conventional methods of clone detection conjoined with their advantages, as presented in this paper, will play a role in the future. We though concede that the assessment of the advantages and disadvantages of conventional code clone detection in comparison with Deep Learning approaches, as well as their incorporation into a possible combined approach, requires further research and analysis. Besides, we are also interested in extending StoneDetector with respect to other programming languages, which have not been in the research focus for code clone detection. Finally, we believe that StoneDetector’s runtime can be improved by further optimizing its implementation as well as examining configurations which trade off accuracy against better runtime, as has been discussed for the Hamming distance.